The length of a rectangle is 3 cm greater than its width. The perimeter is 24 cm. Find the dimensions of the rectangle.
Answer by checkley79(3341)

Perimeter=2L+2W
L=W+2
24=2(W+2)+2W
24=2W+4+2W
24-4=4W
20=4W
20/4=W
5=W ANS. FOR THE WIDTH
L=5+2=7 ANS. FOR THE LENGTH.
PROOF:
24=2*7+2*5
24=14+10
24=24